OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4).

OpenStudy (anonymous):

hint : m1=m2

OpenStudy (anonymous):

Slopes are equal

OpenStudy (anonymous):

Slope = -a/b = 4/3

OpenStudy (anonymous):

nw use y-y1=m(x-x1)